# The Medium Is Not the Message: Deconfounding Text Embeddings via Linear Concept Erasure

### v2

@@ Abstract
Embedding-based similarity metrics between text sequences can be influenced not just by the content dimensions we most care about, but can also be biased by spurious attributes like the text’s source or language. These document confounders cause problems for many applications, but especially those that need to pool texts from different corpora. This paper shows that a debiasing algorithm that removes information about observed confounders from the encoder representations substantially reduces these biases at a minimal computational cost. Document similarity and clustering metrics improve across every embedding variant and task we evaluate—often dramatically. Interestingly, performance on out-of-distribution benchmarks is not impacted, indicating that the embeddings are not otherwise degraded [^1] We publish our code at https://github.com/y-fn/deconfounding-text-embeddings..
@@
Deconfounding Text Embeddings via Linear Concept Erasure
Yu Fan1 Yang Tian2 Shauli Ravfogel3 Mrinmaya Sachan1Elliott Ash1thanks: Equal supervision.Alexander Hoyle [^1] footnotemark: 1 1ETH Zurich 2University of Zurich 3Center for Data Science, New York University
@@ 1 Introduction
Suppose a political scientist is studying U.S. political discourse. They have access to data from two sources: Twitter posts from senators and summaries of congressional bills. A natural first step in data exploration is to first embed the texts (e.g., with a sentence transformer; Reimers and Gurevych 2019) and then cluster them (e.g., with $k$-means). As it turns out, some clusters will overwhelmingly contain items from one source or the other, because systematic differences between sources dominate the distances underpinning $k$-means (Fig. 1A)Correpondence to Yu Fan: yufan@ethz.ch; Yang Tian: yang.tian@uzh.ch; Shauli Ravfogel: shauli.ravfogel@gmail.com; Mrinmaya Sachan: msachan@ethz.ch; Elliott Ash: ashe@ethz.ch; Alexander Hoyle: holyea@ethz.ch.
@@
Figure 1: Clustering text embeddings from disparate sources (here, U.S. congressional bill summaries and senators’ tweets) can produce clusters where one source dominates (Panel A). Using linear erasure to remove the source information produces more evenly balanced clusters that maintain semantic coherence (Panel B; sampled items relate to immigration). Four random clusters of $k$-means shown ($k$=25), trained on a combined 5,000 samples from each dataset.
@@
Text embeddings are generated by pretrained models, and are able to capture topical, semantic, stylistic, multilingual, syntactic, and other information about the embedded text. Generally, models are trained with the goal of “making semantically similar sentences close in vector space” Reimers and Gurevych (2019). But pushing for this goal means that spurious correlations between attributes—e.g., domain and topic—can lead models to learn unwanted relationships. Per Thompson and Mimno (2018): “collections are often constructed by combining documents from multiple sources, [so the] most prominent patterns in a collection simply repeat the known structure of the corpus.” [^2] Their focus is on bag-of-words topic models rather than text embeddings, so their vocabulary-based approach does not translate to our setting. It would therefore seem useful to remove unwanted information from the representations.
@@
Adjusting embeddings to remove confounding information is exactly what we do in this work. Adapting the algorithm from Belrose et al. (2023) for linear concept erasure, we remove embedding subspaces that are predictive of the confounding variables that can bias measures of document distance. In the above example from U.S. politics, we residualize out the source information (Twitter or bills), producing adjusted embeddings for which similarity metrics load on the semantic content rather than the source (Fig. 1B). As another practical example, in a multilingual corpus, we residualize out the subspace that is predictive of language, leading to document distance metrics that are driven by content, rather than language.
@@
We show through extensive tests that the adjusted embeddings perform significantly better for clustering and similarity search. For example, in a multilingual document search setting, Recall@1 increases from 0.175 to 0.826. Intriguingly, there is also no reduction in performance when using the adjusted embeddings on unseen datasets and tasks from a standard retrieval benchmark Muennighoff et al. (2023); Enevoldsen et al. (2024), suggesting erasure does not harm embedding quality.
The approach is computationally inexpensive, involving only linear transformations on pretrained embeddings. Further, it can be used to adjust the embeddings for documents that don’t have labels for the confounders. As a result, the method is particularly useful for applied work, for example in computational social science research. In sum, we:
@@
  • Formally show how erasure removes confounding information from document similarities (§2);
@@
  • Construct a benchmark of paired data designed to measure the effects of confounding attributes on embedding performance (§3);
@@
  • Evaluate a varied set of embedding methods, establishing that observable features, like a text’s source, can harm the utility of text embeddings in applied settings (§4);
@@
  • Demonstrate that applying a linear erasure algorithm to remove observed confounders can mitigate such issues—sometimes dramatically—without impacting other aspects of performance (§5).
@@ 2 Background
Many downstream tasks—nearest-neighbor search, clustering, retrieval, topic discovery—reduce to assessing how “close” two documents are in an embedding space. A good distance metric should rank pairs by semantic relatedness rather than by superficial attributes such as author, language, or publication venue. In practice, however, pretrained embedding models often encode these incidental signals because they appear frequently during training and help optimize their self-supervised objectives. When such signals correlate with content, distance measures become biased, undermining empirical conclusions drawn from them.
@@
Sentence-level embeddings place semantically close documents near one another in a vector space (Kiros et al., 2015; Conneau et al., 2017; Cer et al., 2018; Reimers and Gurevych, 2019). Modern systems start with a transformer encoder trained on masked-language modeling, then refine it on hundreds of millions of contrastive pairs drawn from many corpora (Reimers and Gurevych, 2019). This recipe powers state-of-the-art results in retrieval (Asai et al., 2021; Thakur et al., 2021; Zhang et al., 2023), clustering (Aggarwal and Zhai, 2012), and classification (Maas et al., 2011).
Contrastive batches often contain items from a single source so the model can focus on internal semantics (Nussbaum et al., 2024). A side-effect is that separate sources may occupy separate regions of the space, especially when cross-source positives are scarce. Multilingual models face a similar issue: even when trained with translation pairs (Wang et al., 2024), large amounts of monolingual data still push languages apart.
Notwithstanding efforts to make contrastive pairs comparable, the resulting embeddings still encode confounding information. Platform-specific jargon and style can be pivotal. Language can proxy for topic or geography. For authors or outlets, stylistic markers linked to gender or ideology become shortcuts for similarity. Because these attributes correlate with content, they act as observed confounders in distance-based analyses.
@@
More formally, consider pairs of documents $d_{0}$ and $d_{1}$ with unit-norm embeddings $\smash{\bm{x_{0}},\bm{x_{1}}\in\mathbb{R}^{d}}$, $\lVert\bm{x_{i}}\rVert=1$. Assume a linear decomposition for the embedding:
$\bm{x}_{i}\;=\;\bm{B}_{z}\bm{z}_{i}\;+\;\bm{B}_{c}\bm{c}_{i}\;+\;\bm{B}_{u}\bm{u}_{i}\;+\;\varepsilon_{i},$ (1)
where $\bm{z}_{i}$ captures the semantic content of interest (e.g. topic); $\bm{c}_{i}$ collects observed confounders (source, language, author traits); $\bm{u}_{i}$ collects unobserved confounders; $\bm{B}_{z}$,$\bm{B}_{c}$, $\bm{B}_{u}$ are loading matrices; and $\varepsilon_{i}$ is mean-zero noise uncorrelated with $(\bm{z}_{i},\bm{c}_{i},\bm{u}_{i})$; we also assume these factors are zero-mean and have zero covariance with each other.
Similarity is measured with the dot product:
$Y_{01}\;=\;Y(\bm{x}_{0},\bm{x}_{1})\;=\;\bm{x}_{0}^{\!\top}\bm{x}_{1}.$ (2)
Taking expectations and using (1) gives
$\mathbb{E}[Y_{01}]\;=\;\bm{z}_{0}^{\!\top}\bm{\Gamma}_{z}\bm{z}_{1}\;+\;\bm{c}_{0}^{\!\top}\bm{\Gamma}_{c}\bm{c}_{1}\;+\;\bm{u}_{0}^{\!\top}\bm{\Gamma}_{u}\bm{u}_{1}.$ (3)
where $\bm{\Gamma}_{k}$  =  $\bm{B}_{k}^{\!\top}\bm{B}_{k}$. Only the first term reflects the semantic proximity we care about; the others bias any analysis based on $Y_{01}$.
@@
Early debiasing work on word vectors identified a “bias direction” (e.g. gender) and removed its projection (Bolukbasi et al., 2016). Subsequent studies showed that the removed signal was still recoverable (Gonen and Goldberg, 2019), prompting stronger linear methods such as INLP (Ravfogel et al., 2020), LACE (Ravfogel et al., 2022), and LEACE (Belrose et al., 2023). These approaches search for an affine map that destroys all linear correlation with a protected attribute while moving points as little as possible.
An important special case of these kinds of concept erasure is linear concept erasure, where the goal is to prevent linear adversaries from predicting the information we aim to remove. This is usually achieved in the form of a projection matrix that neutralizes a subspace that is associated with the concept $C$. Following Ravfogel et al. (2022), Belrose et al. (2023) derived sufficient and necessary conditions for achieving linear guardedness (Ravfogel et al., 2023), a situation where no linear classifier can recover the concept $C$ and achieve a loss lower than that of a trivial predictor that always predicts the majority class. Specifically, they derive a linear projection matrix P* such that:
$\bm{P}*$ $=\arg\min_{\mathbf{P}\in\mathbb{R}^{d\times d}}\mathbb{E}\left[||\bm{P}\mathbf{x}-\mathbf{x}||\right]$ (4)
subject to $\text{subject to Cov}(\bm{P}\textbf{x},C)=0.$ (5)
@@
$\bm{P}=\bm{I}\;-\;\bm{W}^{\dagger}\!(\bm{W}\bm{\Sigma}_{XC})(\bm{W}\bm{\Sigma}_{XC})^{\dagger}\bm{W}$ (6)
where $\bm{W}=\bm{\Sigma}_{XX}^{-1/2}$ is a whitening matrix, and
$\bm{\Sigma}_{XC}=\operatorname{Cov}(X,C),\bm{\Sigma}_{XX}=\operatorname{Cov}(X),\bm{\mu}=\mathbb{E}[X].$
This condition is proved to be sufficient and necessary for achieving linear guardedness, i.e., the inability of any linear classifier to recover the attribute $C$ from the embeddings.
In other words, Eq. (6) and $\bm{b}=\bm{\mu}-\bm{P}\bm{\mu}$ give the unique affine map that removes all linear correlation with the observed confounder $C$ while altering the embeddings as little as possible.
For any (possibly unlabeled) document, the adjusted embedding $\tilde{\bm{x}}_{i}\;=\;P\bm{x}_{i}+\bm{b}$ has $\operatorname{Cov}(\tilde{\bm{x}},C)=0$ while minimizing the distance between $\bm{x}$ and $\tilde{\bm{x}}$. Applying the LEACE map $\tilde{\bm{x}}_{i}=\bm{P}\bm{x}_{i}+\bm{b}$ to the structural decomposition in (1) gives
$\tilde{X}_{i}$ $=\bm{P}\bigl{(}\bm{B}_{Z}\bm{z}_{i}+\bm{B}_{C}\bm{c}_{i}+\bm{B}_{U}\bm{u}_{i}+\varepsilon_{i}\bigr{)}+\bm{b}$
$=\bm{B}_{Z}\tilde{\bm{z}}_{i}+\bm{B}_{U}\tilde{\bm{u}}_{i}$ (7)
where the middle term vanishes; since $\operatorname{Cov}(\bm{P}\bm{x},C)=0$, $\bm{B}_{C}=0$. In turn, the estimand for the document similarity
$\tilde{Y}_{01}\;=\;\tilde{\bm{x}}_{0}^{\!\top}\tilde{\bm{x}}_{1}\;=\;\tilde{\bm{z}}_{0}^{\!\top}\bm{\Gamma_{Z}}\tilde{\bm{z}}_{1}\;+\;\tilde{\bm{u}}_{0}^{\!\top}\bm{\Gamma}_{U}\tilde{\bm{u}}_{1},$ (8)
is also purged of $C$. Note, however, that $\tilde{\bm{z}}=\bm{P}\bm{z}$ may not be equal to $\bm{z}$, depending on the intensity and nature of the dependence between $\bm{z}$ and $C$. So the LEACE algorithm might also add bias to similarity metrics through its adjustment of $\bm{z}$. Further, the (adjusted) unobserved confounder $\tilde{\bm{u}}$ remains, and it is unclear how the deconfounding by LEACE would either increase or reduce bias from $\bm{u}$.
3 Experimental Setup
Our evaluation settings are designed to approximate real-world use cases involving datasets from multiple corpora. They are divided into two groups, category-level and event-level data, both aiming to measure the same thing: the extent to which documents that share a common label have similar embeddings.
3.1 Category-level Data

### v3

# The Medium Is Not the Message: Deconfounding Document Embeddings via Linear Concept Erasure

@@ Abstract
Embedding-based similarity metrics between text sequences can be influenced not just by the content dimensions we most care about, but can also be biased by spurious attributes like the text’s source or language. These document confounders cause problems for many applications, but especially those that need to pool texts from different corpora. This paper shows that a debiasing algorithm that removes information about observed confounders from the encoder representations substantially reduces these biases at a minimal computational cost. Document similarity and clustering metrics improve across every embedding variant and task we evaluate—often dramatically. Interestingly, performance on out-of-distribution benchmarks is not impacted, indicating that the embeddings are not otherwise degraded. [^1] Code and data available at https://github.com/y-fn/deconfounding-embeddings.
@@
Deconfounding Document Embeddings via Linear Concept Erasure
Yu Fan1 Yang Tian2 Shauli Ravfogel3 yufan@ethz.ch yang.tian@uzh.ch shauli.ravfogel@gmail.com Mrinmaya Sachan1 Elliott Ash1thanks: Equal supervision. Alexander Hoyle [^1] footnotemark: 1 msachan@ethz.ch ashe@ethz.ch hoylea@ai.ethz.ch 1ETH Zurich  2University of Zurich  3Center for Data Science, New York University
@@ 1 Introduction
Suppose a political scientist is studying U.S. political discourse. They have access to two data sources: Twitter posts from senators and summaries of congressional bills. A natural first step in data exploration is to embed the texts (e.g., with a sentence transformer; Reimers and Gurevych 2019) and then cluster them (e.g., with $k$-means). However, some clusters will predominantly contain items from one source or the other, because systematic differences between sources dominate the distances that $k$-means relies on (Fig.˜1A).
@@
Figure 1: Clustering text embeddings from disparate sources (here, U.S. congressional bill summaries and senators’ tweets) can produce clusters where one source dominates (Panel A). Using linear erasure to remove the source information produces more evenly balanced clusters that maintain semantic coherence (Panel B; sampled items relate to immigration). Four random clusters of $k$-means shown (kk=25), trained on a combined 5,000 samples from each dataset.
@@
Text embeddings, generated by pretrained models, capture a wide range of information about text, including topical, semantic, stylistic, multilingual, and syntactic features. These models are typically trained with the goal of “making semantically similar sentences close in vector space” Reimers and Gurevych (2019). However, this objective can cause spurious correlations—such as between domain and topic—to be encoded as unintended relationships. As Thompson and Mimno (2018) observe: “collections are often constructed by combining documents from multiple sources, [so the] most prominent patterns in a collection simply repeat the known structure of the corpus.” [^2] Their analysis focuses on bag-of-words topic models rather than text embeddings, so their vocabulary-based approach does not translate to our setting. It therefore would seem useful to remove unwanted information from text representations.
@@
Indeed, adjusting embeddings to remove confounding information is exactly what we do in this work. Adapting the algorithm from Belrose et al. (2023) for linear concept erasure, we remove embedding subspaces that are predictive of the confounding variables, which can bias measures of document distance. In the above example from U.S. politics, we residualize out the source information (Twitter or bills), producing adjusted embeddings for which similarity metrics load on the semantic content rather than the source (Fig.˜1B). As another practical example, in a multilingual corpus, we residualize out the subspace that is predictive of language, leading to document distance metrics that are driven by content, rather than language.
@@
Extensive tests show that the adjusted embeddings perform significantly better for clustering and similarity search. For example, in a multilingual document search setting, Recall@1 increases from 0.18 to 0.83. Importantly (and surprisingly), there is also no reduction in performance when using the adjusted embeddings on unseen datasets and tasks from a standard retrieval benchmark Muennighoff et al. (2023); Enevoldsen et al. (2024), suggesting erasure does not harm embedding quality.
The approach is computationally inexpensive, involving only linear transformations on pretrained embeddings. To support practical use, we release a wrapper that streamlines encoding, fitting the deconfounder, and adjusting embeddings. In addition, we provide several evaluation datasets labeled both with confounders (language and source) and semantic content (e.g., topic). In sum, we:
@@
  • Formally demonstrate how erasure removes confounding information from document similarities (§2);
@@
  • Construct a benchmark of paired data to measure the impact of confounding attributes on embedding performance (§3);
@@
  • Evaluate a diverse set of embedding methods, showing that observable features such a text’s source can reduce the utility of text embeddings in applied settings (§4);
@@
  • Show that applying a linear erasure algorithm to remove observed confounders can effectively mitigate such issues—sometimes dramatically—without degrading other aspects of performance (§5).
Figure 2: Performance of the multilingual E5-base model before and after erasure. Top: category-level results from the Comparative Agendas Project (metric: purity). Bottom: event-level results from SCOTUS and multilingual Swiss case summaries (metric: recall@1). Detailed results for other models and datasets in Tables˜2, 4, 10 and 6.
@@ 2 Background
Many downstream tasks such as nearest-neighbor search, clustering, retrieval, topic discovery etc. rely on assessing how “close” two documents are in an embedding space. An effective distance metric should rank pairs by semantic relatedness rather than by superficial attributes like author, language, or publication venue. In practice, however, pretrained embedding models often encode these incidental signals, since they occur frequently during training and help optimize self-supervised objectives. When such signals correlate with content, distance measures become biased and can undermine empirical conclusions.
@@
Sentence-level embeddings position semantically similar documents close to each other in a vector space (Kiros et al., 2015; Conneau et al., 2017; Cer et al., 2018; Reimers and Gurevych, 2019). Modern systems typically begin with a transformer encoder pretrained on masked-language modeling, then refine it on hundreds of millions of contrastive pairs drawn from diverse corpora (Reimers and Gurevych, 2019). This approach underpins state-of-the-art performance in retrieval (Asai et al., 2021; Thakur et al., 2021; Zhang et al., 2023), clustering (Aggarwal and Zhai, 2012), and classification (Maas et al., 2011).
Contrastive batches are often drawn from a single source, enabling the model to focus on internal semantics (Nussbaum et al., 2024). A side effect is that different sources may occupy distinct regions of the embedding space, especially when cross-source positives are scarce. Multilingual models exhibit a similar problem: even when trained with translation pairs (Wang et al., 2024), large amounts of monolingual data tends to push languages apart.
Despite attempts to construct comparable contrastive pairs, the resulting embeddings still encode confounding information. Platform-specific jargon and style can dominate representations. In addition, language itself may serve as a proxy for topic or geography. For authors and outlets, stylistic markers linked to gender or ideology can act as shortcuts for similarity. Because these attributes correlate with content, they function as observed confounders in distance-based analyses.
@@
More formally, let $\bm{X}\in\mathbb{R}^{d}$ denote the embedding of a random document. For a particular document $d_{i}$, we write its realization as $\bm{x}_{i}$ and assume $\lVert\bm{x}_{i}\rVert=1$. In general, we use bold uppercase letters (e.g., $\bm{X}$) for random variables and lowercase letters ($\bm{x}_{i}$) for their realizations. Erased variables use a tilde ($\tilde{\mathrm{B}}$). Assume a linear decomposition for the embedding:
[…]
Similarity is measured with the dot product, where $x_{0}$ and $x_{1}$ are two IID draws of $\bm{X}$:
$y_{01}\;=\;\bm{x}_{0}^{\!\top}\bm{x}_{1}.$ (2)
Under the model expressed in (1), expanding this dot product creates multiple entangled factors:
$y_{01}\;=\;\bm{z}_{0}^{\!\top}\Gamma_{zz}\bm{z}_{1}\;+\;\bm{z}_{0}^{\!\top}\Gamma_{zc}\bm{c}_{1}$ (3)
$\;+\;\bm{z}_{0}^{\!\top}\Gamma_{zu}\bm{u}_{1}\;+\;\bm{c}_{0}^{\!\top}\Gamma_{zc}\bm{z}_{1}\;+\;\ldots,$ (4)
where $\Gamma_{jk}$  =  $B_{j}^{\!\top}B_{k}$. Only the first term reflects the semantic proximity we care about; the others bias any analysis based on $y_{01}$.
@@
Concept erasure techniques aim to remove a targeted feature (e.g. gender) from an embedding. This is typically achieved by projecting out the corresponding subspace, thereby reducing bias and enabling analysis of model behavior without that feature (Ravfogel et al., 2022; Belrose et al., 2023).
Early debiasing work on word vectors identified a “bias direction” (e.g. race) and removed its projection (Bolukbasi et al., 2016). Later studies showed that the removed signal remained recoverable (Gonen and Goldberg, 2019), motivating stronger linear methods such as Iterative Null-space Projection (INLP, Ravfogel et al., 2020), Linear Adversarial Concept Erasure (LACE, Ravfogel et al., 2022), and LEAst-squares Concept Erasure (LEACE, Belrose et al., 2023). These approaches seek an affine transformation that eliminate all linear correlation with the protected attribute while altering the representations as little as possible.
An important special case of these kinds of concept erasure is linear concept erasure, where the goal is to prevent linear adversaries from predicting the information we aim to remove. This is usually achieved in the form of a projection matrix that neutralizes a subspace that is associated with the concept $\bm{C}$. Following Ravfogel et al. (2022), Belrose et al. (2023) derived sufficient and necessary conditions for achieving linear guardedness (Ravfogel et al., 2023), a situation where no linear classifier can recover the concept $\bm{C}$ and achieve a loss lower than that of a trivial predictor that always predicts the majority class. Specifically, they derive a linear projection matrix $\bm{P}^{\ast}$ such that:
$\mathrm{P}^{\ast}$ $=\arg\min_{\mathrm{P}\in\mathbb{R}^{d\times d}}\mathbb{E}\left[\|\mathrm{P}\bm{X}-\bm{X}\|^{2}\right]$ (5)
$\text{subject to Cov}(\mathrm{P}\bm{X},\bm{C})=0.$ (6)
@@
$\mathrm{P}^{\ast}=\mathrm{I}\;-\;\mathrm{W}^{\dagger}\!(\mathrm{W}\mathrm{\Sigma}_{XC})(\mathrm{W}\mathrm{\Sigma}_{XC})^{\dagger}\mathrm{W}$ (7)
where $\mathrm{W}=\mathrm{\Sigma}_{XX}^{-1/2}$ is a whitening matrix, $\mathrm{W}^{\dagger}$ the pseudoinverse of $\mathrm{W}$, and
$\mathrm{\Sigma}_{XC}=\operatorname{Cov}(\bm{X},\bm{C}),\mathrm{\Sigma}_{XX}=\operatorname{Cov}(\bm{X}),\bm{\mu}=\mathbb{E}[\bm{X}].$
This condition is proved to be sufficient and necessary for achieving linear guardedness, i.e., the inability of any linear classifier to recover the attribute $\bm{C}$ from the embeddings. In other words, Eq. (7) together with $\bm{b}=\bm{\mu}-\mathrm{P}\bm{\mu}$ defines the unique affine map that removes all linear correlations with the observed confounder $\bm{C}$ while modifying the embeddings as little as possible.
For any document, the adjusted embedding is $\tilde{\bm{x}}_{i}=\mathrm{P}\bm{x}_{i}+\bm{b}$. Applying this LEACE map to the realization of the structural decomposition in (1):
$\tilde{\bm{x}}_{i}$ $=P\bigl(\mathrm{B}_{z}\bm{z}_{i}+\mathrm{B}_{c}\bm{c}_{i}+\mathrm{B}_{u}\bm{u}_{i}+\bm{\varepsilon}_{i}\bigr)+\bm{b}$
$=\tilde{\mathrm{B}}_{z}\bm{z}_{i}+\tilde{\mathrm{B}}_{u}\bm{u}_{i}+\mathrm{P}\bm{\varepsilon}_{i},$ (8)
where $\tilde{\mathrm{B}}_{j}=\mathrm{P}\mathrm{B}_{j}$. Note that the middle term vanishes, following from the constraint $\operatorname{Cov}(\tilde{\bm{X}},\bm{C})=0$, which ensures that the space spanned by $\mathrm{B}_{c}$ is removed. In turn, the estimand for the document similarity,
$\tilde{y}_{01}\;=\;\tilde{\bm{x}}_{0}^{\!\top}\tilde{\bm{x}}_{1}\;=\;\bm{z}_{0}^{\!\top}\tilde{\mathrm{\Gamma}}_{zz}\bm{z}_{1}\;+\;\bm{z}_{0}^{\!\top}\tilde{\mathrm{\Gamma}}_{zu}\bm{u}_{1}$
$\;+\;\bm{u}_{0}^{\!\top}\tilde{\mathrm{\Gamma}}_{uz}\bm{z}_{1}\;+\;\bm{u}_{0}^{\!\top}\tilde{\mathrm{\Gamma}}_{uu}\bm{u}_{1}+P\bm{\varepsilon}_{i},$ (9)
is also purged of $\bm{C}$ (furthermore, in expectation it does not include the cross terms, as they are uncorrelated under our assumptions). Note, however, that the projection alters the geometry of the remaining components. $\tilde{\bm{x}}$ is now based on $\mathrm{P}\mathrm{B}_{z}\bm{z}$, which may not be equal to $\mathrm{B}_{z}\bm{z}$, depending on the intensity and nature of the dependence between $\bm{z}$ and $\bm{C}$. So the LEACE algorithm might also add bias to similarity metrics through its adjustment of this term. Further, the (adjusted) unobserved confounder $\bm{u}$ remains, and it is unclear how the deconfounding by LEACE would either increase or reduce bias from $\bm{u}$.
@@
3 Experimental Setup
Our evaluation settings are designed to approximate real-world use cases and involve datasets from multiple corpora. They are divided into two groups, category-level and event-level data, both aiming to measure the same thing: the extent to which documents that share a common label have similar embeddings.
3.1 Category-level Data